OpenStudy (anonymous):

Find the lateral area of a cone with a radius of 7 ft. and a slant height of 13 ft. Use 3.14 for π and round to the nearest tenth. (1 point) 439.6 ft2 324.5 ft2 571.5 ft2 285.7 ft2 2. Find the surface area of a square pyramid with a base length of 24 cm and a height of 16 cm. (1 point) 1056 cm2 1536 cm2 816 cm2 1344 cm2

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

@TuringTest you think you can help me?(:

OpenStudy (turingtest):

what is this a test on?

OpenStudy (anonymous):

its not a test, its a questions on a worsheet for volume and radius.

OpenStudy (turingtest):

good, in that case I can tell you that the formula for a cone is \[\frac13\pi r^2h\]where h is the vertical height of the cone. you only are given the slant height and radius, but you can get the vertical height from the pythagorean theorem

OpenStudy (turingtest):

formula for the *Volume* of a cone... anyhow, you can get the height h like so....

OpenStudy (turingtest):

r^2+h^2=x^2 solve for h, plug that into the formula for the volume of a cone

OpenStudy (anonymous):

alright,  im a very confused person right now